OpenStudy (anonymous):

A relatedTo A = {(1,1), (1,2), (2,2), (3,4), (4,1), (4,4)} This statement is AntiSymmetric; But, I don't understand why. Help?

OpenStudy (anonymous):

Antisymmetric  For every element x or y in the set, for every case where (x,y) and (y,x) are both in the relation, x = y.

OpenStudy (freckles):

so for it to by antisymmetric we must have if R(a,b) and R(b,a) then a=b. The relation is vacuously true since you do not have (2,1) or (4,3)...You have the elements to satisfy the hypothesis.

OpenStudy (freckles):

oops didn't see the rest of the set of (1,4)

OpenStudy (freckles):

The relation being antisymmetric is vacuously true is what I meant

OpenStudy (freckles):

remember the following true table: p q if p then q T T T T F F F T T F F T ---- we have the third/fourth line going on

OpenStudy (anonymous):

I am not seeing it. If we have (1,2) should we not also have (2,1). That would kinda of seem like x=y, then.

OpenStudy (freckles):

we don't have (2,1) in your set is what I'm saying

OpenStudy (anonymous):

and, I am still trying to figure out how you drew up that truth table. I understand the truth table. But, where did it come from? What are the p's and q's?!

OpenStudy (freckles):

we can't fulfill the hypothesis of your statement it makes no different if the conclusion is true or not

OpenStudy (freckles):

p is the hypothesis of your statement (the if part) q is the conclusion of your statement (the then part) p is (a,b) in R and (b,a) in R q is a=b

OpenStudy (freckles):

p is false in your case since you not have (a,b) in R and (b,a) in R

OpenStudy (freckles):

so it makes no difference if q is false or true according to the true table the statement is true

OpenStudy (anonymous):

Hold on, let me read over this and see if I can't soak it in.

OpenStudy (anonymous):

@freckles Okay, so in order for this to not be anti symmetric we need (x,y)(y,x) for every element in the set. Not just one?

OpenStudy (freckles):

A={(1,1),(1,2),(2,2),(3,4),(4,1),(4,4),(2,1)} this one would not be antisymmetric since we (1,2) and (2,1) we do have 1 does not equal 2

OpenStudy (freckles):

I just added a point to your original set to make it not antisymmetric

OpenStudy (anonymous):

Wouldn't (1,2)(2,1) mean 1=2?

OpenStudy (freckles):

do you really think 1 is 2 though ?

OpenStudy (freckles):

1=2 is not true

OpenStudy (freckles):

1 is a different number then 2

OpenStudy (anonymous):

Antisymmetric For every element x or y in the set, for every case where (x,y) and (y,x) are both in the relation, x = y.

OpenStudy (anonymous):

x is different than y but says x=y here

OpenStudy (freckles):

right so it is NOT antisymmetric since we have to have 1=2 for it to be antisymmetric

OpenStudy (freckles):

but 1=2 is totally false

OpenStudy (anonymous):

for 1=2 we need (1,2)(2,1) no?

OpenStudy (anonymous):

how else can you make a=b (which i have written 1=2 because we were talking about the elements 1 and 2 (1,2)(2,1))

OpenStudy (anonymous):

Do you see where my confusion is?

OpenStudy (freckles):

(1,2) and (2,1) is in R is true but 1=2 is not true

OpenStudy (freckles):

why are you forcing it to be antisymmetric when it is not ?

OpenStudy (freckles):

if it doesn't satisfy the definition it just doesn't satisfy you cannot make 1 be 2 ever

OpenStudy (anonymous):

I can see it's not antisymmetric!

OpenStudy (freckles):

ok good

OpenStudy (anonymous):

For every element x or y in the set, for every case where (x,y) and (y,x) are both in the relation, x = y. Okay, this is the definition from my book. What I am trying to clarify is that for (1,2)(2,1), by this definition, it seems to me to say that 1=2. Because, in the first pair x=1 and y=2 in the second x=2 and y=1. The end of the definition says that when this happens x=y.

OpenStudy (freckles):

A = {(1,1), (1,2), (2,2), (3,4),(4,1), (4,4)} A is antisymmetric since we cannot satisfy the hypothesis of antisymmetric (think true table) B={(1,1),(1,2),(2,2),(3,4),(4,1),(4,4)} B is not antisymmetric because we were able to satisfy the hypothesis of antisymmetric but we ere not able to satisfy the conclusion of the definition (think true table again if you need)

OpenStudy (anonymous):

I can see from what you said it's not antisymetric. That's no problem. There is no 2,1 and that's good enough for me. But, I don't understnad the last part of that definition.

OpenStudy (freckles):

the then part says a has to be equal to b but we have 1 is not 2 if 1 was 2 then yes it would by antisymmetric

OpenStudy (freckles):

but as I said we know 1 is not 2

OpenStudy (freckles):

if they were the same they would have a difference of 0 2-1=1 <--not 0

OpenStudy (freckles):

2 is greater than 1 actually

OpenStudy (freckles):

so it cannot be equal to

OpenStudy (freckles):

are you understanding the definition of antisymmetric is: (x,y) in R and (y,x) in R then x=y you understand we do not have this conclusion is true for that set B, right since 2 is not 1?

OpenStudy (anonymous):

Okay, so I think I may have figured out where I was getting confused. (x,y) and (y,x) means x=y because both y and x have two values

OpenStudy (freckles):

we plug into the definition and got True for hyp and false for conlusion which makes the statement false

OpenStudy (anonymous):

if they had 1 value they could not be equivalent. I concede that 1>2

OpenStudy (anonymous):

erm

OpenStudy (anonymous):

2>1 :P

OpenStudy (anonymous):

Hyp = (1,2) = T and (2,1) doesn't exist so False, yes? This doesn't imply a true conclusion.

OpenStudy (freckles):

ok I was trying to think of a set that is antisymmetric that has the hyp and conclusion satisfy (and I know this example is going to be a dumb example but it is all I can of so far: C={(x,1),(1,x) | x=1} is antisymmetric why? well it satisfies the hyp: we do have (x,1) and (1,x) is in the relation we also have the conclusion since x=1 :p

OpenStudy (freckles):

right I think I get what you are saying... B={(1,2),(2,1)} is not antisymmetric because when we try to plug into the definition which is: if (x,y) and (y,x) are in the relation then x=y we get if (2,1) and (1,2) are in the relation (which is true) then 1=2 (which is false) truth table tells us then this statement is false so it is not antisymmetric

OpenStudy (anonymous):

Yes, I think you can see what my issue was now. I was trying to plug this into the definition; but, I was only giving x and y 1 value. They, have two though. It might have made more sense if I had said "x has values 1 and 2. y has values 1 and 2. therefore, x=y."

OpenStudy (freckles):

so based on what you are you think E={(1,2),(2,2)} is antisymmetric or not antisymmetric? also how about D={(x,y) in IntegerXInteger|y=x-1 or x=y-1}

OpenStudy (anonymous):

E is antisymmetric. I'm still trying to figure out D. I'm assuming it would be antisymmetric also, because there is only 1 ordered pair.

OpenStudy (freckles):

That is right on E lol I'm lost on D too... I kinda made it up... well 5 and 6 integers which means and 5=6-1 so (6,5) is in D (because of the y=x-1 thing) is (5,6) also in D? well 5=6-1 from the x=y-1 thing so it does seem to be antisymmetric -- thinking... a little more

OpenStudy (anonymous):

I'm thinking, too. Am trying to, at least :P

OpenStudy (freckles):

ok let's look at the lines y=x-1 and x=y-1 solving the second equation for y gives y=x+1 f(x)=x-1 and g(x)=x+1 are inverses since f(g(x))=f(x+1)=x+1-1=x and g(f(x))=g(x-1)=x-1+1=x so since f and g are inverses that means since (a,b) is in D then (b,a) is in D that is for example: since (7,8) is in D then (8,7) is in D but 7 is not 8 err so I think is not antisymmetric

OpenStudy (freckles):

so I don't know why I said it was antisymmetric earlier I was looking at (5,6) and (6,5) being in D but 6 is not 5

OpenStudy (freckles):

that should have been the end of it

OpenStudy (anonymous):

6>5!

OpenStudy (anonymous):

:D

OpenStudy (freckles):

lol yes :p

OpenStudy (freckles):

that was interesting question slightly harder

OpenStudy (anonymous):

I've gota get some sleep Dr. Freckles. I'm sure I'll have more questions for you tomorrow if you're around. And, I'll leave this question up so I can re-read your brainteaser tomorrow. Believe it or not, this stuff does interest me. I'm just not very good at it-- yet...

OpenStudy (freckles):

i can understand the definitions being confusing at first

OpenStudy (freckles):

anyways goodnight

OpenStudy (anonymous):

Good night :)